OpenStudy (anonymous):

Simplify: (8-4x)/(2x-4)

OpenStudy (anonymous):

need help fast

OpenStudy (anonymous):

factor out a-4 from the top and 2 from the bottom -4(x-2)/2(x-2) cancel the two with the four and the stuff in parentheses also cancels = -2

OpenStudy (anonymous):

factor out a-4 from the top and 2 from the bottom -4(x-2)/2(x-2) cancel the two with the four and the stuff in parentheses also cancels = -2

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

-2

jhonyy9 (jhonyy9):

-4(x-2) ------- = -4/2 = -2 2(x-2)

jhonyy9 (jhonyy9):

- hope that you will can understanding it